OpenStudy (anonymous):

QUESTION: WORD PROBLEM - FARM ANIMALS: I AM A FARMER AND NEED TO PURCHASE SOME FARM ANIMALS YOU ARE GOING TO SELL ME 100 FARM ANIMALS FOR $100 DOLLARS. (NO MORE , NO LESS) COWS COST $5.00 EACH PIGS COST $1.00 EACH CHICKENS COST $0.5 EACH TOTAL $100.00 FOR 100 FARM ANIMALS I NEED TO KNOW HOW MANY COWS, HOW MANY CHICKENS, HOW MANY PIGS I CAN GET FOR MY $100.00 BUT REMEMBER, I MUST HAVE 100 FARM ANIMALS ALL TOTAL. I CANNOT HAVE 99. AND I CANNOT HAVE 101. PLEASE NOTE THE ANSWER IS NOT 100 PIGS.

OpenStudy (anonymous):

let cow be x pigs be y chickens be z x+y+z= 100 5x + y + 0.5z = 100

OpenStudy (anonymous):

That does not tell me how many cows, chickens and pigs. Are you saying the answer is creating an equation for it?

OpenStudy (anonymous):

Creating equation is first step. Nest step is to solve it. can you help me in that

OpenStudy (anonymous):

I thought the answer was simply 100 pigs == $100.00 my teacher said it was incorrect.

OpenStudy (anonymous):

@StudyGurl14 need help here

OpenStudy (studygurl14):

Looks like you have the equation set up perfect @sangya21 :) What do you need help with?

OpenStudy (anonymous):

I am not able to solve it. elimination method will give 0 constant

OpenStudy (studygurl14):

Hold on a sec...

OpenStudy (studygurl14):

x+y+z= 100 5x + y + 0.5z = 100 -x - y - z = -100 5x + y + 0.5z = 100 4x + 0.5z = 0 x+y+z= 100 y = 100 - x - z 5x + (100 - x - z) + 0.5z = 100 4x - 0.5z = 0 4x + 0.5z = 0 4x - 0.5z = 0 8x = 0 x = 0

OpenStudy (studygurl14):

4(0) + 0.5z = 0 z = 0 Okay...yeah this doesn't make sense...the answer HAS to be 100 pigs...

OpenStudy (studygurl14):

I suppose the answer could also be 100/0.5 = 200 chickens...

OpenStudy (studygurl14):

or...100/5 = 20 cows

OpenStudy (anonymous):

LOL that is what I turned in and it was wrong. It can only be 100 animals for $100 dollars no more no less.

OpenStudy (studygurl14):

yea, this problem is stupid. I give up lol

OpenStudy (studygurl14):

it is impossible...

OpenStudy (anonymous):

LOL! THANKS FOR TRYING!

OpenStudy (studygurl14):

You're welcome. :) sorry i couldn't help. :(

OpenStudy (anonymous):

When I get the answer in class I will return and give it to you!

OpenStudy (studygurl14):

Please do! I would like to know the answer to this one...

OpenStudy (anonymous):

Ok, signing off now!

OpenStudy (anonymous):

Yeah thanks